# Privacy risk in GeoData: A survey

### v1

@@ Abstract.
With the ubiquitous use of location-based services, large-scale individual-level location data has been widely collected through location-awareness devices. The exposure of location data constitutes a significant privacy risk to users as it can lead to de-anonymisation, the inference of sensitive information, and even physical threats. Geoprivacy concerns arise on the issues of user identity de-anonymisation and location exposure. In this survey, we analyse different geomasking techniques that have been proposed to protect the privacy of individuals in geodata. We present a taxonomy to characterise these techniques along different dimensions, and conduct a survey of geomasking techniques. We then highlight shortcomings of current techniques and discuss avenues for future research.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXX
@@ 1. Introduction
Data privacy is a problem that exists in any data that is about, caused by, or initiated by people (Westin, 1968). Organisations in many business domains increasingly produce large databases with millions of records, which might contain detailed and sensitive information about people, such as customers, patients, taxpayers, or travellers (Drechsler and Hu, 2021). Consequently, concern about the privacy of these data has become a top priority. A violation of the privacy of an individual is the use of personal information other than the norm allows. Thus, data privacy can be viewed as a guideline for how data should be collected or handled, based on its sensitivity and importance. Although contextual integrity clearly defines how a privacy violation could occur in a dataset, it provides no protection mechanism other than policy and regulations (Zandbergen, 2014).
@@
Geo spatial information (we refereed to such data as geodata) at the most detailed level (coordinates) has been recognised as important information in various domains such as epidemiology, medicine, and social science (Fronterrè et al., 2018). The amount of high-resolution spatial data has been enormously growing due to the enclosing of positioning capabilities in mobile, wearable, and global positioning systems. Geo spatial data can be linked with other information to create geocoded data, such as mapping individual’s address details into geographic coordinates. As more and more data is available about individuals, geocoded data enables the accurate identification of individuals’ locations (Sharad and Danezis, 2013; De Montjoye et al., 2013; Unnikrishnan and Naini, 2013).
@@
As geocoded data are becoming available at increasingly finer resolutions, attacker capabilities in linking geocoded data to specific individuals or groups continue to grow. Knowing an individual’s location increases the risk of reidentification by reverse geocoding and can seriously violate the individual’s privacy (Brownstein et al., 2006a; Kounadi et al., 2013). For instance, when publishing a data point, coordinates can be simply calculated and connected with an address. Thus, the address attribute is supposed to be processed as confidential information because it can reveal an individual and, thereby, ought to be deleted before publishing a dataset.
@@
The growing need for protecting discrete geodata results from technological developments and the need of releasing data online. Hence, over the past decade there have been many papers published that proposed different methods to preserve the individual’s privacy, such as geomasking techniques (Zandbergen, 2014). Geomasking techniques are meant to secure confidentiality when publishing geodata while preserving geographic detail to enable a precise spatial analysis of the dataset (Kwan et al., 2004a; Gambs et al., 2014; Wang et al., 2022). These techniques aim to obfuscate the real locations of an individual. For example, using geomasking the address of an individual does not have to be eliminated but can be transferred to a different position thus avoiding the deletion of the addresses from a published dataset results in a less accurate data utility.
Geomasking techniques aim to obfuscate the real locations of an individual. However, there is no generally recommended nor approved geomasking techniques. Each geomasking method has disadvantages and advantages. Moreover, no geomasking method currently applicable is able to serve a thorough solution to protect location privacy. Thus, an in-depth analysis of the capabilities of each geomasking technique is required to better utilise these geomasking techniques in real applications.
Contribution: In this paper we conduct a comprehensive survey on geomasking techniques. To guide our survey of geomasking techniques, we propose a taxonomy that we used to review geomasking technique under different dimensions in this article. At a high level, this taxonomy differentiates between geomasking techniques in terms of the way how each technique can be used to anonymise geo data, different evaluation measures that can be used to evaluate these technique, and practical aspects that are important when implementing and experimenting these techniques. The goal of this proposed taxonomy is to allow data custodians to easily navigate the large list of existing mechanisms and identify the mechanisms that could be fit-for-purpose for their needs. We acknowledge that other taxonomies could have been equally adopted and that some of these alternatives may be more useful for other readers of this survey (e.g., scientists looking for knowledge gaps in the area).
Comparison to other works: Several surveys have studied the geomasking techniques. Boulos et al. (Kamel Boulos et al., 2022) provides a comprehensive overview of location privacy issues and geoprivacy-preserving methods within the realm of public health interventions and health research with aggregate geographic data. The survey delves into critical challenges related to location privacy, particularly in the context of public health crises like the COVID-19 pandemic.
Maike et al. (Steffen et al., 2023) systematically reviewed the existing literature on data protection strategies for geocoded data and classifies them into aggregation, geographic masking, and synthetic data. The survey also discusses various tools for assessing the risk and utility of protected geocoded data. (Kim et al., 2021) investigated how do different geomasking methods impact perceived disclosure risk. The findings reveal that perceived disclosure risk increases with the amount of locational information on a map. The study suggests that some geomasking methods such as aggregation methods could contribute to decreased perceived disclosure risk.
Some studies have explored the applicability of geomasking techniques in specific domains. Nowbakht et al. (Nowbakht et al., 2022) investigated the challenge of balancing confidentiality and spatial pattern preservation in sharing agricultural data. (Tiwari et al., 2023) focuses on the ethical use of geographical information systems (GIS) data, specifically discussing geomasking for locational privacy. (Hu, 2019) focuses on Bayesian methods for estimating disclosure risks in synthetic data, emphasising their construction and application. Iyer et al. (Iyer et al., 2023) categorises geomasking algorithms into aggregation, affine transformation, and random perturbation, offering a comprehensive exploration of each method’s advantages and drawbacks. The study also delves into the evaluation criteria for geomasking procedures, emphasising the importance of analytical accuracy and privacy protection. The thesis(Kekana, 2020) focuses on evaluating four masking methods—grid masking (GM), random perturbation (RP), weight random perturbation (WRP), and Voronoi masking (VM)—with the goal of presenting the sleeping locations of the homeless without compromising their spatial confidence.
@@
(Broen et al., 2021) investigates the impact of various data perturbation methods on both key spatial statistics and patient privacy. The study investigated distinct perturbation approaches ranging from non-aggregating techniques like random perturbation, and Voronoi masking, to aggregating perturbations such as grid line masking and grid center masking. (Wang et al., 2022) conducted a comprehensive evaluation of eight geomasking methods using simulated geospatial data with diverse spatial patterns to assess their effectiveness in privacy protection and analytical accuracy. The recent thesis by Redlich (Redlich, 2022) presents a comprehensive evaluation of geomasking methods designed to balance individual privacy protection with the preservation of spatial information utility. This work provides valuable insights into geographic masking methods’ strengths, limitations, and privacy implications, forming a basis for informed decision-making in spatial data analysis. However, the authors did not analyse the geomasking techniques under the dimensions that are important for practical applications but only viewed from a risk and utility perspective.
Further, different studies (Dupre et al., 2022; Löffler et al., 2017; Seidl et al., 2018; Gao et al., 2019) conducted comprehensive explorations of challenges associated with the increasing availability of geocoded data and the growing capabilities of attackers to link such data to specific individuals or groups. Overall, these studies highlights the need for further research to compare the effectiveness of different masking methods, determine optimal displacement levels for confidentiality, and explore alternatives to geographic masking for protecting individual-level geodata.
@@ 2.1. Different types of Geodata
Geo data consists information that describes objects, events or other features with a location on or near the surface of the earth. Geo data is information recorded in conjunction with a geographic indicator of some type. There are two primary forms of geo data: vector data and raster data. Vector data is data in which points, lines and polygons represent features such as properties, cities, roads, mountains and bodies of water. For example, a visual representation using vector data might include houses represented by points, roads represented by lines and entire towns represented by polygons. Raster data is pixelated or gridded cells which are identified according to row and column. Raster data creates imagery that’s substantially more complex, such as photographs and satellite images.
Apart from vector and raster data, location information can be come in the forms of census data, mobility data, drawn images, and social media data. In census data can contain data that tied to specific geographical areas and thus statistical analysis can be conducted on data attached to different geographical locations. Mobility data provides information regarding locations and movements of individuals. Individuals will generate mobility data about their movements through tagging locations on social media, using apps which collect location data such as mapping apps, through interacting with wifi beacons, or in some cases from bluetooth, GPS or mobile phone records.
Mobility data can usually be captured as GPS location coordinates and also based on the connectives an individual’s mobile make with mobile cell towers. Such data can also be used to perform geospatial analysis. Drawn images such as CAD images providing information of building or other structures providing geographical information. Finally, social media data can also contain data about people movements as they post information regarding the places that a person has visited.
@@ 2.2. Different Privacy attacks on Geodata
Geodata privacy refers to a set of principles and practices put in place to protect the confidentiality of individuals,groups, communities or organisations about whom there is geographically explicit data collected. Geocoded data are becoming available at increasingly finer resolutions, while attacker capabilities in linking geocoded data to specific individuals or groups continue to grow. The number of identifying variables necessary for a data breach to occur is reduced dramatically when geographic location is known at a precise level, especially where population density is low. The development of geocoded data sources presents new concerns for data confidentiality breaches, as traditional disclosure avoidance mechanisms may ignore spatial characteristics. In this section we explore different privacy issues for geodata.
@@ 2.2.1. Reverse Geocoding
Geocoding refers to the process of assigning geographic coordinates to addresses or place names. However, the release of geographic information at the individual level can breach privacy and confidentiality. For example, publishing the street address of an individual makes it possible to look up the associated name(s) in directories and property databases. Publishing a location as coordinates means that these can be plotted on a map and then associated with an address. Publishing a map in paper or digital form also means that the locations can be associated with an address.
Reverse geocoding occurs when the address details for a location published is determined. Address details can then be a powerful key for linking the data associated with those coordinates to specific and identifiable information from other sources. For example, in an application domain of census data, the locations of the respondents can be inferred and reverse geocoded if a census area has low-value cells leading to possible disclosure. Reverse geocoding can lead to re-identification because the street address can then be associated with one or several individuals using common directories. Figure 1 illustrates an example of geocoding and reverse geocoding processes of addresses (Kounadi et al., 2013).
Several research works have shown that re-identification of individual addresses using reverse geocoding. In 2006, Brownstein and others (Brownstein et al., 2006a) were able to correctly identify more than 70% of the addresses in a published map using manual reverse geocoding techniques in GIS. The same authors showed similar accuracy in identifying addresses can be achieved using semi-automated reverse geocoding based on image analysis (Brownstein et al., 2006b). In realistic examples, a group of researches from US correctly identified most of the original residence of mortality locations of Hurricane Katrina using a published map in a local newspaper (Curtis et al., 2006). Further, another study of crime incidents in Vienna, Austria showed that individuals could be accurately re-identified using online reverse geocoding and online address and telephone directories.
@@ 2.2.2. Geocoding of administratively-masked data
Helderop et al. (Helderop et al., 2023) recently proposed a geocoding method for unmasking administratively-masked data where specific identifying features, like addresses, are obscured to protect privacy. The proposed approach utilises a medoid-based technique, aiming to minimise spatial uncertainty associated with the masking process. The method utilises a master address point dataset to generate a list of candidate addresses for each masked observation, and a minimum bounding polygon is drawn around these candidates. The medoid, representing the central-most point, is then identified within this polygon and assigned as the geocoded address. The approach outperforms commercial geocoding software, providing higher accuracy and a spatial confidence metric for each result.
2.2.3. Inferencing attack
An inference attack corresponds to a process by which an adversary that has access to some data related to individuals (and potentially some auxiliary information) tries to deduce new personal information that was not explicitly present in the original data. For geographical data, the inferencing attacks have been studied for a long time (Sharad and Danezis, 2013; De Montjoye et al., 2013; Unnikrishnan and Naini, 2013). The most common inferencing attack type applied on geo data is de-anonymising attacks (Gambs et al., 2014). The de-anonymising attack aims to re-identify individuals using their anonymised location data.
Gambs et al. (Gambs et al., 2014) proposed a de-anonymising attack where an adversary tries to infer the identity of a particular individual using their mobility traces. More precisely, the proposed attack can re-identify individuals with a high success rate whose movements are contained in an anonymous dataset provided that the adversary can used as background information some mobility traces of the same individuals that he has been able to observe during the training phase.
Wang et al. (Wang et al., 2016) proposed a new model to profile users’ spatio-temporal mobility behaviours and use the model to launch the de-anonymisation attack. The authors used User Hidden Markov model to profile the user movements and used an algorithm to rank the users based on the probability that each individual matches with a given profile.
A recent work by Eshun and Palmieri (Eshun and Palmieri, 2022) proposed two de-anonymisation techniques using hidden markov models. In their attacks both the temporal and spatial influences on user mobility trajectories are considered for making prediction on users movemement. These predictions are used to create a profile for each individual and the re-identification of individuals based on these user profiles are performed based on a divergence measure. An experimental evaluation on real-world mobility traces showed that their proposed attack can achieve over 80% of accuracy in re-identifying individuals.

### v2

@@ Abstract.
With the ubiquitous use of location-based services, large-scale individual-level location data has been widely collected through location-awareness devices. The widespread exposure of such location data poses significant privacy risks to users, as it can lead to re-identification, the inference of sensitive information, and even physical threats. In this survey, we analyse different geomasking techniques proposed to protect individuals’ privacy in geodata. We propose a taxonomy to characterise these techniques across various dimensions. We then highlight the shortcomings of current techniques and discuss avenues for future research. Our proposed taxonomy serves as a practical resource for data custodians, offering them a means to navigate the extensive array of existing privacy mechanisms and to identify those that align most effectively with their specific requirements.
@@
doi: ccs: Security and privacy Data anonymization and sanitizationccs: Security and privacy Pseudonymity, anonymity and untraceabilityccs: Information systems Geographic information systemsccs: Information systems Global positioning systemsccs: Information systems Location based servicesccs: General and reference Surveys and overviews
@@ 1. Introduction
Data privacy refers to individuals’ protection and control over their personal data, particularly concerning its collection, processing, storage, and sharing by organisations (Houser and Voss, 2018; Rustad and Koenig, 2019). It is crucial for any data that pertains to, is caused by, or initiated by people. Organisations across various business sectors increasingly produce large databases with millions of records, potentially containing detailed and sensitive information about individuals, such as customers, patients, taxpayers, or travellers (Drechsler and Hu, 2021). Consequently, safeguarding the privacy of such data has become a paramount concern.
Privacy violation occurs when personal information is utilised beyond permissible norms. Thus, the concept of data privacy serves as a guiding principle dictating the collection and handling of data based on its sensitivity and significance. However, while contextual integrity elucidates how privacy breaches can manifest within a dataset by analysing adherence to social norms specific to different contexts, it lacks protective mechanisms beyond policies and regulation (Nissenbaum, 2004).
@@
Geospatial data (referred to herein as geodata) is recognised as crucial information across various domains, including epidemiology, medicine, and social science (Fronterrè et al., 2018). At its most detailed level, this data comprises coordinates pinpointing specific locations. For example, geodata can be linked with other information to create geocoded data, which maps individuals’ address details into geographic coordinates. The amount of high-resolution spatial data has been growing enormously due to the integration of positioning capabilities into mobile, wearable, and global positioning systems. With the increasing availability of data about individuals, geodata enables the precise identification of individuals’ locations (Sharad and Danezis, 2013; De Montjoye et al., 2013; Unnikrishnan and Naini, 2013).
@@
As geodata become increasingly accessible at finer resolutions, attacker capabilities in linking geodata to specific individuals or groups continue to grow. Knowing an individual’s location increases the risk of reidentification through reverse geocoding and can seriously violate the individual’s privacy (Brownstein et al., 2006a; Kounadi et al., 2013). For instance, when publishing a data point, coordinates can be simply calculated and connected with an address. Therefore, the address attribute should be treated as confidential information because it can reveal an individual.
@@
The growing need for protecting discrete geographic data results from technological developments and the need to release data online. Hence, numerous studies in the past decade have proposed various approaches to preserve sensitive information, such as geomasking techniques (Zandbergen, 2014). Geomasking techniques are meant to preserve privacy when publishing geodata while maintaining geographic detail to enable precise spatial analysis of the dataset (Kwan et al., 2004a; Gambs et al., 2014; Wang et al., 2022). These techniques aim to obfuscate the real locations of an individual. For instance, by employing geomasking techniques, the address of an individual does not have to be eliminated but can be transferred to a different position, which would result in less accurate data utility but higher privacy.
Despite the various techniques proposed, there are no generally recommended nor approved geomasking techniques. Each presents its unique set of advantages and disadvantages. Furthermore, none of the current masking methods offer a comprehensive solution to protect location privacy. Therefore, a thorough analysis of the capabilities of each technique is imperative to harness them effectively in practical applications. Moreover, ongoing research and refinement are necessary to enhance the efficacy of geomasking strategies in preserving privacy.
@@
Comparison to other works: Several surveys have studied the geomasking techniques. Steffen et al. (Steffen et al., 2023) systematically review the existing literature on data protection strategies for geocoded data and classify them into aggregation, geographic masking, and synthetic data. The survey also discusses various tools for assessing the risk and utility of protected geocoded data. The recent thesis by Redlich (Redlich, 2022) presents a comprehensive evaluation of geomasking methods designed to balance individual privacy protection with the preservation of spatial information utility. This work provides valuable insights into geographic masking methods’ strengths, limitations, and privacy implications, forming a basis for informed decision-making in spatial data analysis. Wang et al. (Wang et al., 2022) conduct a comprehensive evaluation of eight geomasking methods using simulated geospatial data with diverse spatial patterns to assess their effectiveness in privacy protection and analytical accuracy.
Some studies have explored the applicability of geomasking techniques in specific domains. Kamel Boulos et al. (Kamel Boulos et al., 2022) provide a comprehensive overview of the challenges surrounding location privacy and techniques for preserving geoprivacy within the realm of public health interventions and health research that utilise detailed individual geographic data. (Iyer et al., 2023; Broen et al., 2021; Zandbergen, 2014) explore the use of geomasking techniques for perturbing spatial statistics and patient privacy in epidemiologic research and health analytics. Nowbakht et al. (Nowbakht et al., 2022) investigate the challenge of balancing confidentiality and spatial pattern preservation in sharing agricultural data. The thesis by Kekana (Kekana, 2020) focuses on evaluating geomasking methods on presenting the sleeping locations of the homeless without compromising their spatial confidence. Gao et al.(Gao et al., 2019) examine the performance of random and Gaussian perturbation methods for Twitter users sharing their geotagged tweets.
Further, different studies address a specific concern within the geoprivacy domain (Solymosi et al., 2023). Seidl et al. (Seidl et al., 2018) explore the assessment of the risk of false identification in geomasking techniques. Dupre et al. (Dupre et al., 2022) propose to manage privacy risks associated with geocoded data through geospatial disclosure avoidance techniques while Tiwari et al. (Tiwari et al., 2023) investigate the ethical utilisation of geographic information system (GIS) data using geomaking techniques. Kim et al. (Kim et al., 2021) investigate how do different geomasking methods impact perceived disclosure risk and Hu (Hu, 2019) explore the utilisation of Bayesian methods for estimating disclosure risks in synthetic data.
Overall, all the proposed studies discuss the different geomasking techniques and challenges associated with the increasing availability of geocoded data. To the best of our knowledge, there has not yet been a thorough survey of geomasking techniques and classify these techniques under various aspects. This article aims to fill that gap.
Contribution: We present an extensive survey of geomasking techniques. We introduce a taxonomy that reviews geomasking methods across various dimensions, namely privacy techniques, evaluation measures, and practical considerations. At a high level, this taxonomy differentiates between geomasking techniques based on how each technique can anonymise geodata, the diverse evaluation measures applicable to them, and practical aspects. These practical aspects encompass the experimental datasets, purpose-built prototypes designed specifically for implementing these techniques, and the wide range of application domains where these techniques find relevance. The goal of this taxonomy is to facilitate data custodians in navigating the extensive array of existing mechanisms, aiding them in identifying those best suited to their needs. Furthermore, this survey endeavors to highlight any existing gaps in these techniques, offering valuable insights for future researchers to explore and refine. While our survey provides guidance across various aspects, we acknowledge that other taxonomies proposed by different studies could have been equally adopted. Some of these taxonomies may offer more detailed insights for specific inquiries (e.g., scientists interested in exploring knowledge gaps in the field). Additionally, we offer an overview of advanced large language models (LLMs) specifically designed to tackle geospatial data analytics. While these LLMs may not directly address privacy issues, they provide a valuable framework for researchers to explore and integrate privacy considerations, enabling them to address complex geoprivacy issues within their methodologies.
@@ 2.1. Different types of Geodata
Geodata comprises information describing objects, events, or features with a location on Earth, recorded alongside a geographic indicator. It primarily exists in two forms: vector data and raster data. Vector data delineates features such as properties, cities, roads, mountains, and bodies of water, represented by points, lines, and polygons. For instance, a visual representation using vector data might include houses represented by points, roads by lines, and entire towns by polygons. Conversely, raster data consists of pixelated or gridded cells, creating complex imagery like photographs and satellite images.
Besides vector and raster data, location information manifests in various other forms, including census data, mobility data, drawn images, and social media data. Census data provides insights into specific geographical areas, facilitating statistical analysis across different locations. Mobility data offers insights into individuals’ locations and movements, which can be collected through tagging locations on social media, using location-tracking apps, interacting with WiFi beacons, or through Bluetooth, Global Positioning System (GPS), or mobile phone records. This data is valuable for conducting geospatial analysis. Drawn images, such as Computer-aided design (CAD) images, provide information about buildings or other structures, offering geographical insights. Furthermore, social media data can include information about people’s movements as they post about their visited places.
@@ 2.2. Different Privacy attacks on Geodata
Geodata privacy refers to a set of principles and practices put in place to protect the privacy of individuals, groups, communities or organisations whose data is represented in geographically explicit formats (Kamel Boulos et al., 2022). With geodata increasingly available at finer resolutions and attackers continuously advancing their capabilities in linking geodata to specific individuals or groups, the risk of a privacy breach escalates, especially when precise geographic locations are known, particularly in areas with low population density. The emergence of various geodata sources introduces new concerns for privacy breaches, as traditional disclosure avoidance mechanisms may overlook spatial characteristics (Bertino et al., 2008; Solymosi et al., 2023). In this section, we explore a range of privacy attacks/breaches targeting geodata.
@@ 2.2.1. Reverse Geocoding
Geocoding is the process of assigning geographic coordinates to addresses or place names, enabling spatial data analysis and visualisation. However, releasing geographic information at the individual level can pose significant privacy risks. For instance, disclosing a street address can allow individuals to be easily identified through directories and property databases. Similarly, publishing coordinates enables pinpoint locations on maps, potentially revealing sensitive information. Whether in print or digital format, maps containing such data can inadvertently disclose addresses and compromise privacy.
Reverse geocoding occurs when the address details for a published location are determined. These address details serve as powerful keys for linking the data associated with those coordinates to specific and identifiable information from other sources. For example, in the context of census data, the respondents’ locations can be inferred and reverse geocoded if a census area has cells with low values, potentially leading to inadvertent disclosure. Reverse geocoding can facilitate the re-identification of individuals, as it allows the conversion of geographic coordinates into street addresses. Once a street address is obtained, it becomes possible to use common directories to associate that address with one or several individuals. Figure 1 illustrates an example of geocoding and reverse geocoding processes for addresses (Kounadi et al., 2013).
Several research works have shown the re-identification of individual addresses using reverse geocoding. In 2006, Brownstein and others (Brownstein et al., 2006a) were able to correctly identify more than 70% of the addresses in a published map using manual reverse geocoding techniques in GIS. The same authors showed similar accuracy in identifying addresses can be achieved using semi-automated reverse geocoding based on image analysis (Brownstein et al., 2006b). In realistic examples, a group of researchers from the US correctly identified most of the original residents of mortality locations of Hurricane Katrina using a published map in a local newspaper (Curtis et al., 2006). Further, another study of crime incidents in Vienna, Austria, showed that individuals could be accurately re-identified using online reverse geocoding and online address and telephone directories (Kounadi et al., 2013).
@@ 2.2.2. Geocoding of administratively-masked data
Helderop et al. (Helderop et al., 2023) propose a geocoding method for unmasking administratively-masked data where specific identifying features, like addresses, are obscured to protect privacy. The proposed approach utilises a medoid-based technique, aiming to minimise spatial uncertainty associated with the masking process. The method utilises a master address point dataset to generate a list of candidate addresses for each masked observation, and a minimum bounding polygon is drawn around these candidates. The medoid, representing the central-most point, is then identified within this polygon and assigned as the geocoded address. The approach outperforms commercial geocoding software, providing higher accuracy and a spatial confidence metric for each result.
2.2.3. Inference attack
An inference attack corresponds to a process by which an adversary that has access to some data related to individuals (and potentially some auxiliary information) tries to deduce new personal information that was not explicitly present in the original data. For geographical data, the inference attacks have been studied for a long time (Sharad and Danezis, 2013; De Montjoye et al., 2013; Unnikrishnan and Naini, 2013). The most common inference attack type applied on geodata is de-anonymising attacks (Gambs et al., 2014). The de-anonymising attack aims to re-identify individuals using their anonymised location data.
Gambs et al. (Gambs et al., 2014) introduce a de-anonymising attack strategy, where an adversary tries to infer the identity of a specific individual through analysis of their mobility traces. More precisely, this attack method demonstrates a high success rate in re-identifying individuals whose movements are documented in an anonymous dataset. This success is contingent upon the adversary having access to mobility traces of the same individuals, observed during a training phase, which they can employ as background information.
Wang et al. (Wang et al., 2016) propose a new model to profile users’ spatiotemporal mobility behaviours and use the model to launch the de-anonymisation attack. The authors use the User Hidden Markov Model to profile the user movements and used an algorithm to rank the users based on the probability that each individual matches with a given profile.
A recent work by Eshun and Palmieri (Eshun and Palmieri, 2022) proposes two de-anonymisation techniques using Hidden Markov Models. In their attacks, both the temporal and spatial influences on user mobility trajectories are considered for predicting user movements. These predictions are then used to create a profile for each individual, and the re-identification of individuals based on these user profiles is achieved through a divergence measure. An experimental evaluation on real-world mobility traces shows that their proposed attack can achieve over 80% of accuracy in re-identifying individuals.